solving by substitution method. 4x-5y= -35
4x+39=y

Answer by stanbon(75887)

solving by substitution method.
4x-5y= -35
4x+39= y
----------------
Substitute for "y" and solve for "x":
4x-5(4x+39) = -35
4x-20x-195 = -35
----
-16x = 160
x = -10
---------------
Solve for "y":
y = 4x+39
y = 4(-10)+39
y = -40+39
y = -1
===================
Cheers,
Stan H.